OpenStudy (anonymous):

What is the equation of the line passing through (8, –6) and perpendicular to the line y = x – 2 in slope-intercept form? Answer y = –x – 2 y = x + 2 y = –x + 2 y = x – 2

OpenStudy (akshay_budhkar):

y=-x+2.. Do u want me to explain?

OpenStudy (anonymous):

yes please

OpenStudy (akshay_budhkar):

ok we know the gerenal equation of a line is y=mx+c where m is the slope. Now if two lines are perpendicular the we have the product of their slopes to be -1. therefore in this case the slope of the line perpendicular to the original one is -1

OpenStudy (akshay_budhkar):

so the equation of the line is y=-x+c now it is given that it passes through ( 8,-6) so we subsititute the values to get the value of c

OpenStudy (akshay_budhkar):

um am i clear?